OpenStudy (anonymous):

What is (x+y)^(2/3)?

ganeshie8 (ganeshie8):

\(\large (x+y)^{\frac{2}{3}} \)

ganeshie8 (ganeshie8):

\(\large [(x+y)^2]^{\frac{1}{3}} \)

ganeshie8 (ganeshie8):

\(\large \sqrt[3]{(x+y)^2} \)

ganeshie8 (ganeshie8):

is that wat u looking for ?

OpenStudy (anonymous):

Can you expand it?

ganeshie8 (ganeshie8):

oh binomial expansion !

OpenStudy (anonymous):

To clarify, how would you expand (x+y)^(2/3)?

OpenStudy (anonymous):

To solve Lim h--> 0 (f(a+h)-f(a))/h for f(x) = x^(2/3)?

ganeshie8 (ganeshie8):

ahh that should be easy, \(\large \lim \limits_{h \to 0} \frac{(x+h)^\frac{2}{3}-x^{\frac{2}{3}}}{h}\)

ganeshie8 (ganeshie8):

like that ?

OpenStudy (anonymous):

Yes! But then how would you expand (x+h)^(2/3)?

OpenStudy (anonymous):

If you don't use the limit way, how can I prove f(x) is differentiable at every point on (0,1) for the mean value theorem?

OpenStudy (anonymous):

f(x) refers to x^(2/3)

ganeshie8 (ganeshie8):

not sure, sorry :(

OpenStudy (agent0smith):

Yeah i tried and gave up. Wolfram alpha used l'hopitals rule to solve it, idk if there's a way to simplify that.

OpenStudy (anonymous):

I do know l'hopitals but we didn't learn it in class yet so it's not allowed. I think I'll just use agent0smith's way of showing x in (0,1) is valid for f'(x)

OpenStudy (anonymous):

What is the degree of (1/x)-((x+1)/2x)?

OpenStudy (agent0smith):

degree usually refers to polynomials... I guess that would be degree zero if you really needed the degree, since (x+1)/2x = x/2x + 1/2x and x/2x = 2. But i don't see why you'd want the degree of that...

OpenStudy (anonymous):

My teacher wanted us to find it. Sometimes, he asks us questions that have no answers to make us think. Maybe this is one of them? No wonder my math hw is taking so long...

OpenStudy (agent0smith):

I think you'd just say it's not a polynomial. Rational functions don't have a degree.

OpenStudy (anonymous):

Good answer! How long will you be online today? Your math is really good!

OpenStudy (agent0smith):

Well i have a degree in maths and physics and i'm a precalculus/algebra 2/AP physics teacher :D i'll be online a bit longer

OpenStudy (anonymous):

You're teaching so many subjects! And you're def the most patient teacher on this website!

OpenStudy (agent0smith):

yeah i teach four total... a physics class in addition to AP physics.

OpenStudy (anonymous):

wow, that's a lot of grading...

OpenStudy (anonymous):

are you online everyday?

OpenStudy (agent0smith):

Yeah, pretty much every day :)

OpenStudy (anonymous):

That's awesome!